# Herald: A Natural Language Annotated Lean 4 Dataset

### v1

@@ Abstract
Verifiable formal languages like Lean have profoundly impacted mathematical reasoning, particularly through the use of large language models (LLMs) for automated reasoning. A significant challenge in training LLMs for these formal languages is the lack of parallel datasets that align natural language with formal language proofs. To address this challenge, this paper introduces a novel framework for translating the Mathlib4 corpus (a unified library of mathematics in formal language Lean 4) into natural language. Building upon this, we employ a dual augmentation strategy that combines tactic-based and informal-based approaches, leveraging the Lean-jixia system, a Lean 4 analyzer. We present the results of this pipeline on Mathlib4 as Herald (Hierarchy and Retrieval-based Translated Lean Dataset). We also propose the Herald Translator, which is fine-tuned on Herald. Herald translator achieves a 93.2% accuracy (Pass@128) on formalizing statements in the miniF2F-test and a 22.5% accuracy on our internal graduate-level textbook dataset, outperforming InternLM2-Math-Plus-7B (74.0% and 7.5%) and TheoremLlama (50.1% and 4.0%). Furthermore, we propose a section-level translation framework for real-world applications. As a direct application of Herald translator, we have successfully translated a template section in the Stack project, marking a notable progress in the automatic formalization of graduate-level mathematical literature. Our model, along with the datasets, will be open-sourced to the public soon.
@@ 1 Introduction
Based on the Herald dataset, we fine-tuned a model for NL-FL statement translation. To validate the generated formalized statements, we apply both Lean compiler and LLM back-translation checks. Our model achieves 93.2% accuracy on miniF2F-test(Zheng et al., 2022) and 22.5% accuracy on our internal graduate-level textbook dataset, outperforming InternLM2-Math-Plus-7B Ying et al. (2024b)(74.0% and 7.5%) and TheoremLlama (Wang et al., 2024) (55.0% and 4.0%). To demonstrate the model’s effectiveness in autoformalization, we apply the Herald translator to a section of the Stack Project (See Appendix B.1), using DeepSeek-Prover-V1.5 (Xin et al., 2024b) to complete the proofs.
@@
  • We release the Herald translator model, fine-tuned on the Herald dataset, achieving 93.2% accuracy on miniF2F-test and 22.5% on our internal graduate-level dataset, significantly outperforming InternLM2-Math-Plus-7B (74.0% and 7.5%) and TheoremLlama (55.0% and 4.0%).
@@ 3.1.1 Statements Informalization
The first step in our methodology involves extracting essential components from Lean code that encapsulate formal statements. We utilize Lean-Jixia [^1] https://github.com/reaslab/Lean-Jixia, a static analysis tool specifically designed for Lean 4, to extract structured information from Mathlib4. Lean-Jixia parses Lean files to extract key metadata, including theorem declarations, proof structures, and dependency relationships. We select five main components to enhance the FL to NL translation process:
By utilizing this information, LLM can better understand the FL statement and follow the principles of NL statements when it translates the FL statements into NL statements.
@@ 3.1.2 Proofs Informalization
In addition to the structured information provided during the translation of statements, we further extract more detailed components for the translation of proofs. These components include:
  • Formal Statement. The formal statement being proved.
@@ 3.2.2 Augmentation via Mathematics-pretrained LLM
  • Logical Equivalence Rewriting For each formal statement, we generate several logically equivalent informal statements.
    Example we might rewrite “If AA, then BB” as “BB holds given AA”.
  • Abstract Concept Substitution In some cases, lengthy or detailed informal statements can be rephrased using more abstract or higher-level mathematical concepts.
    Example Statement as ‘For given matrix $A$, there exists a matrix $B$, such that $AB=BA=I$’ would be replaced with a more concise, abstract expression or concept like ‘$A$ is non-degenerate’.
  • Omission of Implicit Condition In natural language mathematical discourse, especially in textbooks and research papers, certain conditions are often omitted because they are considered obvious or conventionally understood by the reader.
    Example A theorem might not explicitly mention the requirement that a function be continuous if that is implied by the context.
  • Multi-linguistic Translation We generate additional informal statements by translating the formal statements into Chinese, French and Russian. This results in a set of informal statements that represent how theorems might be articulated in a different linguistic context.
@@ 4 Experiments
We conduct extensive experiments to evaluate the Herald dataset and translator. In Section 4.1, we test the Herald translator on three statement datasets from different topics and compare its performance with other formalization models. Section 4.2 assesses the quality of the Herald dataset through expert inspection, while in Section 4.3, we apply our autoformalization pipeline to a section of the Stack Project and analyze the results.
@@ 4.1.1 Dataset and training
We selected DeepSeek-Prover-Base 7B as our base model due to its extensive training on formal programming languages like Lean, which provides a strong foundation for formal reasoning tasks.
@@ 4.1.2 Validation pipeline
For validation, we adopt the pipeline from the LeanWorkbook project (Ying et al., 2024a), which includes several key steps:
@@ 4.1.3 Result
These datasets were carefully chosen to assess the models’ capabilities across various levels of difficulty and categories of mathematics.
Model miniF2F Extract Theorem College CoT
test valid
TheoremLlama 50.1% 55.6% 4.0% 2.9%
InternLM2-Math-Plus-7B 73.0% 80.1% 7.5% 6.5%
Llama3-instruct 28.2% 31.6% 3.6% 1.8%
Herald 93.2% 96.7% 22.5% 17.1%
@@ 4.2 Summary of case study in Herald quality examination
To assess the mathematical rigor and language style in the Herald dataset, we conducted several case studies. For a comprehensive result, see Appendix C.
@@ 4.3 Autoformalization practice
To demonstrate the capabilities of the Herald translator in auto-formalizing modern mathematics, we selected the Normal Extensions section (Stacks Project Authors, 2018, Tag 09HL) from the Field Theory chapter of the Stacks Project. This section was successfully formalized into a runnable Lean 4 source file, showcasing the effectiveness of our pipeline. For more details on the Stacks Project, see Appendix B.1, and for the formalized output, refer to Appendix B.2.
@@ 5 Conclusions
We release the Herald dataset, which includes 580k valid statements and 44k NL-FL theorem pairs, providing a significant resource for formalization research. Additionally, we fine-tuned a statement formalizer on this dataset, which achieves state-of-the-art accuracy—93.2% on the miniF2F-test and 22.5% on a graduate-level textbook dataset—substantially outperforming existing baselines such as InternLM2-Math-Plus-7B and TheoremLlama. By applying our translator to a section of the Stack Project and leveraging the DeepSeek Prover V1.5, we further demonstrate the practical viability of our approach in auto-formalization.
@@ 6 Acknowledgements
Bin Dong is supported in part by the New Cornerstone Investigator Program.
@@ Appendix A Introduction to formal languages
Lean 4 is an interactive, type-theory based prover. (Moura & Ullrich, 2021) Lean 4 is designed to be highly extensible via its metaprogramming capability, enabling Lean-Jixia to extract important metadata accurately and concisely.
@@
Appendix B Graduate Textbook Auto-formalization
B.1 The Stacks Project
B.2 Formalization of Stacks Project section Normal Extensions
@@
open Polynomial
Normal E K := by
have := h.out
exact Normal.tower_top_of_normal F E K

### v2

@@ Abstract
Verifiable formal languages like Lean have profoundly impacted mathematical reasoning, particularly through the use of large language models (LLMs) for automated reasoning. A significant challenge in training LLMs for these formal languages is the lack of parallel datasets that align natural language with formal language proofs. To address this challenge, this paper introduces a novel framework for translating the Mathlib4 corpus (a unified library of mathematics in formal language Lean 4) into natural language. Building upon this, we employ a dual augmentation strategy that combines tactic-based and informal-based approaches, leveraging the Lean-jixia system, a Lean 4 analyzer. We present the results of this pipeline on Mathlib4 as Herald (Hierarchy and Retrieval-based Translated Lean Dataset). We also propose the Herald Translator, which is fine-tuned on Herald. Herald translator achieves a 93.2% accuracy (Pass@128) on formalizing statements in the miniF2F-test and a 22.5% accuracy on our internal graduate-level textbook dataset, outperforming InternLM2-Math-Plus-7B (74.0% and 7.5%) and TheoremLlama (50.1% and 4.0%). Furthermore, we propose a section-level translation framework for real-world applications. As a direct application of Herald translator, we have successfully translated a template section in the Stack project, marking a notable progress in the automatic formalization of graduate-level mathematical literature. Our model, along with the datasets, are open-sourced to the public. [^1] The model can be found at https://huggingface.co/FrenzyMath/Herald_translator. The code of Herald and LeanSearch can be found at https://github.com/frenzymath/herald_translator.
@@ 1 Introduction
Based on the Herald dataset, we fine-tuned a model for NL-FL statement translation. To validate the generated formalized statements, we apply both Lean compiler and LLM back-translation checks. Our model achieves 96.7% accuracy on miniF2F-test (Zheng et al., 2022) and 23.5% accuracy on our internal graduate-level textbook dataset, outperforming InternLM2-Math-Plus-7B (Ying et al., 2024b) (73.0% and 7.5%) and TheoremLlama (Wang et al., 2024) (55.0% and 4.0%). To demonstrate the model’s effectiveness in autoformalization, we apply the Herald translator to a section of the Stack Project (See Appendix F.1), using DeepSeek-Prover-V1.5 (Xin et al., 2024b) to complete the proofs.
@@
  • We release the Herald translator model, fine-tuned on the Herald dataset, achieving 96.7% accuracy on miniF2F-test and 23.5% on our internal graduate-level dataset, significantly outperforming InternLM2-Math-Plus-7B (73.0% and 7.5%) and TheoremLlama (55.0% and 4.0%).
@@ 3.1.1 Statements Informalization
The first step in our methodology involves extracting essential components from Lean code that encapsulate formal statements. We utilize Lean-Jixia [^2] https://github.com/reaslab/Lean-Jixia, a static analysis tool specifically designed for Lean 4, to extract structured information from Mathlib4. Lean-Jixia parses Lean files to extract key metadata, including theorem declarations, proof structures, and dependency relationships. We select five main components to enhance the FL to NL translation process: head statements, kind, docstrings, neighbor statements, and dependent theorems. These additional structures provide mathematical context and background information for the LLM, thereby reducing the burden on the LLM to infer the mathematical background independently. For an overview of each component, please refer to Appendix B. By utilizing this information, the LLM can better understand the FL statement and follow the principles of NL statements when translating FL statements into NL statements.
@@ 3.1.2 Proofs Informalization
In addition to the structured information provided during the translation of statements, we further extract more detailed components for the translation of proofs and utilize the previous statement translation results. These additional components include: formal statement, informal statement, tactic information, and proof states. For an overview of each component, please refer to Appendix B.
@@ 3.2.2 Augmentation via Mathematics-pretrained LLM
  • Logical Equivalence Rewriting For each formal statement, we generate several logically equivalent informal statements. For example, we might rewrite “If AA, then BB” as “BB holds given AA.”
  • Abstract Concept Substitution In some cases, lengthy or detailed informal statements can be rephrased using more abstract or higher-level mathematical concepts. For example, the statement ‘For given matrix $A$, there exists a matrix $B$, such that $AB=BA=I$’ would be replaced with a more concise, abstract expression or concept like ‘$A$ is non-degenerate’.
  • Omission of Implicit Condition In natural language mathematical discourse, particularly in textbooks and research papers, certain conditions are often omitted because they are considered obvious or conventionally understood by the reader. For example, a theorem might not explicitly mention the requirement that a function be continuous if that is implied by the context.
  • Multi-linguistic Translation We generate additional informal statements by translating the formal statements into Chinese, French and Russian. This results in a set of informal statements that represent how theorems might be articulated in a different language.
@@ 4 Experiments
We conduct extensive experiments to evaluate the Herald dataset and translator. In Section 4.1, we test the Herald translator on three statement datasets from different topics and compare its performance with other formalization models. Section 4.2 assesses the quality of the Herald dataset through expert inspection, while in Section 4.3, we apply our autoformalization pipeline to a section of the Stack Project and analyze the results. Ablation study on the two augmentation methods is presented in Appendix C.
@@ 4.1.1 Dataset and training
We selected DeepSeek-Prover-V1.5-Base 7B as our base model due to its extensive training on formal programming languages like Lean, which provides a strong foundation for formal reasoning tasks.
@@ 4.1.2 Validation pipeline
For validation, we adopt the pipeline from the Lean Workbook (Ying et al., 2024a), which includes several key steps:
@@
For a detailed discussion and case study of the validation pipeline’s effects, especially the NLI check, please refer to Appendix D. We conducted human evaluations on the auto-formalization results of ProofNet (Azerbayev et al., 2022). The results indicate that while the pipeline is not perfectly accurate for evaluating informal-to-formal alignment, it remains a common practice in the field and one of the few practical approaches for large-scale automated evaluation at this time.
@@
Model miniF2F Extract Theorem College CoT
test valid
TheoremLlama 50.1% 55.6% 4.0% 2.9%
InternLM2-Math-Plus-7B 73.0% 80.1% 7.5% 6.5%
Llama3-instruct 28.2% 31.6% 3.6% 1.8%
Herald 96.7% 96.3% 23.5% 16.0%
@@ 4.1.3 Result
These datasets were carefully chosen to assess the models’ capabilities across various levels of difficulty and categories of mathematics. For representative examples in the Extract Theorem and College CoT dataset, please refer to Appendix G.
@@
We present the main experimental results in Table 2. From the table, we observe that the Herald translator achieves a success rate of 96.3% on MiniF2F-Valid, 96.7% on MiniF2F-Test, 23.5% on Extract Theorem, and 16.0% on College COT, demonstrating significant improvement on the MiniF2F dataset and a breakthrough on two new datasets, outperforming all baselines. These results indicate that the Herald translator excels at formalizing natural language statements, particularly in challenging tasks and high-level mathematics. A case study of the Herald translator’s auto-formalization results is presented in Appendix D.
@@ 4.2 Summary of case study in Herald quality examination
To assess the mathematical rigor and language style in the Herald dataset, we conducted several case studies. For a comprehensive result, see Appendix E.
@@ 4.3 Autoformalization practice
To demonstrate the capabilities of the Herald translator in auto-formalizing modern mathematics, we selected the Normal Extensions section (Stacks Project Authors, 2018, Tag 09HL) from the Field Theory chapter of the Stacks Project. This section was successfully formalized into a runnable Lean 4 source file, showcasing the effectiveness of our pipeline. For more details on the Stacks Project, see Appendix F.1, and for the formalized output, refer to Appendix F.2.
@@ 5 Conclusions
We release the Herald dataset, which includes 580k valid statements and 44k NL-FL theorem pairs, providing a significant resource for formalization research. Additionally, we fine-tuned a statement formalizer on this dataset, which achieves state-of-the-art accuracy— 96.7% on the miniF2F-test and 23.5% on a graduate-level textbook dataset—substantially outperforming existing baselines such as InternLM2-Math-Plus-7B and TheoremLlama. By applying our translator to a section of the Stack Project and leveraging the DeepSeek Prover V1.5, we further demonstrate the practical viability of our approach in auto-formalization.
@@ 6 Acknowledgements
This work is supported in part by the New Cornerstone Investigator Program and in part by Ubiquant.
@@ Appendix A Introduction to formal languages
Lean 4 (Moura & Ullrich, 2021) is an interactive, type-theory based prover. Lean 4 is designed to be highly extensible via its metaprogramming capability, enabling Lean-Jixia to extract important metadata accurately and concisely.
@@
Appendix B Structured Information in Lean for RAG
Here, we provide a detailed explanation of each component of the structured information we provide to the LLM, which is used to enhance the FL to NL translation process in data generation. During proof translation, all structured information (excluding kind) for statements is provided, along with additional components to further assist the LLM.
Components for statement translation
Additional components for proof translation
  • Formal Statement. The formal statement being proved.
Appendix C Ablation Study
In this section, we conduct an ablation study on the two augmentation methods. We trained 5 models on different composition of datasets. These training sets share the same ”original” set which denotes the informal-formal pairs chosen from the translated Mathlib4 dataset, and we ablate the augmented datasets by adjusting different ratios between the ”original” set and augmented statements to assess the effectiveness of our methods.
Dataset ratio Dataset size MiniF2F-test ProofNet-valid Extract-Theorem
1:0:0 145k 86.8% 50.0% 15.0%
1:1:1 435k 91.4% 68.3% 17.5%
1:2:0 435k 95.1% 74.7% 19.0%
1:2:1 580k 93.2% 78.6% 20.5%
1:2:1 1160k 96.7% 81.6% 23.5%
Table 3: Performance comparison of Herald translator trained on different size and composition of datasets across various datasets. Dataset ratio represents ‘original : tac_aug : inf_aug‘, original denotes the informal-formal pairs chosen from the translated Mathlib4 dataset, tac_aug refers to the tactic-based augmented statements, and inf_aug corresponds to the informal augmented statements. The penultimate model is trained on NL→FL dataset, and the last model, as our model in the main experiment, is trained on NL→FL and FL→NL datasets.
As shown in table 3, our augmentation methods effectively enhance model performance, particularly on more challenging datasets such as ProofNet. The model trained solely on the original informal-formal pairs (1:0:0) achieves a respectable accuracy of 50.0% on the ProofNet-valid dataset, 86.8% on the MiniF2F-test dataset, and 15.0% on the Extract-Theorem dataset, which underscores the quality of the informalized Mathlib4 dataset.
The introduction of augmented datasets significantly boosts model performance across all datasets. The model trained with a 1:1:1 ratio shows a notable improvement, achieving 68.3% on ProofNet-valid, 91.4% on MiniF2F-test, and 17.5% on Extract-Theorem. The 1:2:0 ratio model, which includes more tactic-based augmented statements, further enhances performance, reaching 74.7% on ProofNet-valid, 95.1% on MiniF2F-test, and 19.0% on Extract-Theorem. The model trained with 1:2:1 ratio on both NL→FL and FL→NL datasets outperforms the model trained with 1:2:1 ratio on only NL→FL datasets, with accuracies of 81.6% on ProofNet-valid, 96.7% on MiniF2F-test, and 23.5% on Extract-Theorem. This result indicates that increasing the size of the training set, while maintaining a balanced augmentation ratio, leads to substantial improvements in model performance, especially on more challenging datasets.
Appendix D Discussion and Case Study for Validation Pipeline
In this section, we discuss the soundness and fairness of our validation pipeline and analyze examples of passing and failing the validation pipeline.
Discussion
It is important to note that even the strictest LLM-based NLI checking introduces false positives. The NLI check consists of two steps: back translation and semantic comparison. During back translation, LLMs are less capable than humans at detecting subtleties in formalized statements, such as filling back opened namespaces and disambiguating overloading notations. This may result in errors in back translation. During semantic comparison, the LLM may fail to decide correctly due to a lack of mathematical understanding and the various possibilities of stating mathematics in neutral language. Consequently, formalized results passing the NLI check may fail to reflect the mathematical meaning of the original natural language statement (i.e., false positives). Therefore, NLI checking is not as reliable as human expert evaluation.
However, while NLI checking is not perfectly accurate for evaluating informal-to-formal alignment, it remains a common practice in the field (e.g., see Ying et al. (2024a)) and one of the few practical approaches for large-scale automated evaluation at this time. It is not practical to apply human evaluation to a dataset with more than 1k formalization results. It is also worth noting that NLI check may introduce false negatives, i.e., results rejected by the NLI check that are actually correct.
For our application of the NLI checking pipeline, the most important goal is to create a fair performance test between different models. As we will see in the following experiments, both the baseline model and our model achieve similar results in terms of the proportion of false-positive cases.
To provide a more illustrative result on false positive cases and fairness mentioned in the above discussion, we conducted an experiment to evaluate the results passing the validation pipeline.
Experiment results
We choose the ProofNet dataset to conduct our experiment. The ProofNet dataset consists of problems drawn from popular undergraduate mathematics textbooks, selected based on the following criteria: self-containment, ease of formalization, and low risk of train-test overlap.
We perform 128 parallel translations of the ProofNet validation set (with 185 natural language statements) using the InternLM2-Math-Plus-7B and Herald model. The translation is considered as validation accepted if any of these passes both the compiler check and the nil check and we only collect the first passed formalization for human evaluation.
The validation-passed auto-formalization outcomes are categorized by human experts into three distinct groups: correct translations, minor errors, and major errors. A correct translation must accurately reflect the mathematical meaning of the natural language statement. In cases where the original statement is ambiguous, we allow the formalized statement to adopt any mathematically valid interpretation. If the formalized result deviates from the original statement but can be easily corrected by altering a single formal definition or restructuring a single logical relation within the hypothesis or statement, it is classified as a minor error. All other cases are classified as major errors. In section D.1, we will present examples of these three categories.
Model Validation Passed Correct Translation Minor Error Major Error
InternLM2-Math-Plus-7B 72 42 12 18
Herald 151 101 24 26
Table 4: Human evaluation results of validation passed statements.
As shown in Table 4, other baseline models also suffer from false positives, with comparable false positive rates, making the comparison fair.
For a more thorough inspection, in subsection D.2, we will present examples that are rejected by the Lean validation step and the NLI validation step. For false negative cases, given the substantial number of translation attempts, it is impractical to manually detect all such cases. Therefore, we will only provide one example in subsection D.2.3.
During the evaluation process, we also identified some potential mathematical inaccuracies or inaccurately truncated problems within the ProofNet dataset, which are detailed in Appendix D.3.
D.1 Results accepted by the validation pipeline
D.1.1 Correct translation
Rudin|exercise_3_20
Informal statement
Suppose $\left\{p_{n}\right\}$ is a Cauchy sequence in a metric space $X$, and some sequence $\left\{p_{nl}\right\}$ converges to a point $p\in X$. Prove that the full sequence $\left\{p_{n}\right\}$ converges to $p$.
Generated formal statement
@@ D.1.1 Correct translation
open Filter
open Topology Filter
theorem full_convergence_of_subsequence {X : Type*} [MetricSpace X] {p : X}
{s : X} (hs : CauchySeq s) (l : ℕ) (hl : Tendsto l atTop atTop)
(hpl : Tendsto (s l) atTop ( p)) : Tendsto s atTop ( p) := sorry
Analysis
This is an example where the Herald translator correctly translated the subtle condition ”some sequence $\left\{p_{nl}\right\}$” in natural language. To be a subsequence, the index $nl$ must tend to infinity, which is reflected in the condition (hl : Tendsto l atTop atTop).
Axler|exercise_5_12
Informal statement
Suppose $T\in\mathcal{L}(V)$ is such that every vector in $V$ is an eigenvector of $T$. Prove that $T$ is a scalar multiple of the identity operator.
Generated formal statement
import Mathlib
open LinearMap
open Submodule
We are to prove that if every vector is an eigenvector,then T is a scalar multiple of identity
theorem isEigenvector_spec (V : Type*) [AddCommGroup V] [Field K] [Module K V]
(T : V →ₗ[K] V) (h : v : V, k : K, k v = T v) : k : K, v : V, k v = T v := by sorry
Analysis
In this example, the formalized condition (h : ∀ v : V, ∃ k : K, k • v = T v) for “every vector in $V$ is an eigenvector” and the formalized result ∃ k : K, ∀ v : V, k • v = T v for “$T$ is a scalar multiple of the identity operator” differ only in the order of quantifiers. The Herald translator handled this difference in the order of quantifiers appropriately.
D.1.2 Minor error
Ireland-Rosen|exercise_18_4
Informal statement
Show that 1729 is the smallest positive integer expressible as the sum of two different integral cubes in two ways.
Generated formal statement
import Mathlib
open Finset
theorem smallest_two_way_sum_of_cubes_is_1729 :
IsLeast {n : | a b c d : ℤ, a c a^3 + b^3 = c^3 + d^3 n = a^3 + b^3} 1729 := sorry
Analysis
In natural language, we treat $1729=a^{3}+b^{3}$ and $1729=b^{3}+a^{3}$ as the same cube decomposition. Therefore, the condition a ≠ c is not sufficient to guarantee two distinct cube decompositions.
Rudin|exercise_1_19
Informal statement
Suppose $a,b\in R^{k}$. Find $c\in R^{k}$ and $r>0$ such that $|x-a|=2|x-b|$ if and only if $|x-c|=r$. Prove that $3c=4b-a$ and $3r=2|b-a|$.
Generated formal statement
import Mathlib
open Topology Filter Set
theorem problem_1_19 (a b : ℝ) : c r : ℝ, |x - a| = 2 * |x - b| |x - c| = r 3 * c = 4 * b - a 3 * r = 2 * |b - a| := by sorry
Analysis
There is a logical flaw in the formalized statement. The formalized statement is equivalent to saying that if $3c=4b-a$ and $3r=2|b-a|$, then $r$ and $c$ satisfy the previous conditions. However, the converse direction, namely that any $r$ and $c$ satisfying the previous conditions must also satisfy $3c=4b-a$ and $3r=2|b-a|$, is not translated.
D.1.3 Major error
Dummit-Foote|exercise_9_4_11
Informal statement
Prove that $x^{2}+y^{2}-1$ is irreducible in $\mathbb{Q}[x,y]$.
Generated formal statement
import Mathlib
open Ideal Ideal.Quotient Finset
theorem irreducible_X_pow_two_add_Y_pow_two_sub_one : Irreducible (X ^ 2 + Y ^ 2 - 1 : Polynomial ℚ) := sorry
Analysis
This is a major error that is difficult to detect. Lean can automatically deduce the type of variables X Y as $\mathbb{Q}$-coefficient polynomials, allowing this statement to pass the Lean check. However, we intended X Y to be the special generators in the ring of polynomials (also in the ring of polynomials in two variables instead of one). This is indeed an incorrect formal statement.
Dummit-Foote|exercise_3_4_11
Informal statement
Prove that if $H$ is a nontrivial normal subgroup of the solvable group $G$ then there is a nontrivial subgroup $A$ of $H$ with $A\unlhd G$ and $A$ abelian.
Generated formal statement
import Mathlib
section exercise_3_4_11